OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

sin^2 x

OpenStudy (p0sitr0n):

u =sinx 7u^2-14u+7=0 u^2-2U+1=0 (u-1)(u-1)=0 root: 1 sinx =1 when x=pi/2, 5pi/2, 9pi/2, ...

OpenStudy (anonymous):

those are the three solutions ? @P0sitr0n

OpenStudy (p0sitr0n):

well, if there is no interval specified, theres infinitely many solutions since sinx is a cyclic function, like it repeats itself